### v1

# Bilevel subsidy-enabled mobility hub network design with perturbed utility coalitional choice-based assignment

Abstract
Urban mobility is undergoing rapid transformation with the emergence of new services. Mobility hubs (MHs) have been proposed as physical-digital convergence points, offering a range of public and private mobility options in close proximity. By supporting Mobility-as-a-Service, these hubs can serve as focal points where travel decisions intersect with operator strategies. We develop a bilevel MH platform design model that treats MHs as control levers. The upper level (platform) maximizes revenue or flow by setting subsidies to incentivize last-mile operators; the lower level captures joint traveler–operator decisions with a link-based Perturbed Utility Route Choice (PURC) assignment, yielding a strictly convex quadratic program. We reformulate the bilevel problem to a single-level program via the KKT conditions of the lower level and solve it with a gap–penalty method and an iterative warm-start scheme that exploits the computationally cheap lower-level problem. Numerical experiments on a toy network and a Long Island Rail Road (LIRR) case (244 nodes, 469 links, 78 ODs) show that the method attains sub–1% optimality gaps in minutes. In the base LIRR case, the model allows policymakers to quantify the social surplus value of a MH, or the value of enabling subsidy or regulating the microtransit operator’s pricing. Comparing link-based subsidies to hub-based subsidies, the latter is computationally more expensive but offers an easier mechanism for comparison and control.
keywords:
Mobility hub , MaaS platform , Perturbed utility route choice , bilevel optimization , subsidies , assignment game
journal: Preprint
1 Introduction
Urban mobility has significantly evolved due to advances in mobile technology and digital platforms. Yet, transportation networks continue to struggle with persistent challenges including congestion, emissions, and unequal accessibility. Despite the proliferation of new mobility services like ride-hailing, car-sharing, and micromobility options, there remain challenges to fully integrate these systems to fundamentally resolve urban mobility issues. Travelers still face inconvenience in seamlessly coordinating multiple modes at transfer locations. Large proportions of commuters may still prefer driving alone if the connections to major public transit stops are inadequate or unreliable.
Emergent cyberphysical mobility hubs (MHs) are a viable response to these challenges. MHs are defined as ”multimodal transport nodes that facilitate intermodal transfers by providing different mobility options in close proximity” (Miramontes et al.,, 2017). While ”transit hubs” have existed for decades as a means to integrate land use and public transport as a type of ”transit-oriented development” (see Weustenenk and Mingardo, (2023)), they differ from MHs in the literature in the last decade, which feature higher levels of both physical and digital integration of mobility services (Arias-Molinares et al.,, 2023), particularly services that feature real-time information. These MHs often involve multiple operators (or even non-mobility-related goods and service providers, such as fueling), necessitating distinct incentives and cooperation mechanisms enabled by integrated digital platforms supporting Mobility-as-a-Service (MaaS). MaaS platforms streamline users’ abilities to plan, book, and pay for various mobility services through a unified digital interface and distribute those revenues between operators, thereby aligning physical infrastructure and digital coordination. As such, MHs form critical components of cyberphysical MaaS ecosystems, acting as strategic nodes for interactions among public entities, private mobility operators, and users.
The deployment of MHs alongside MaaS schemes has been studied globally over the past decade. A notable example is the car-sharing program in Bremen, Germany (Karbaumer and Weltring,, 2025). Car-sharing stations known as “mobil.punkte” are established at easily accessible locations near major public transit stops, with supporting infrastructure such as bike racks and app-based booking systems. Sustainability-focused mobility hub designs have gained more attention with the maturation of electric vehicle (EV) technologies. The eHUBS project, supported by the European Regional Development Fund, investigated the deployment of intermodal hubs focusing on electric-powered modes in six pilot cities across North-West Europe (Amsterdam,, 2025). For a comprehensive review of mobility hub-focused pilot studies worldwide, readers can refer to Arnold et al., (2023).
Designing systems incorporating MHs is no trivial task. Grigolon et al., (2025) reveals that travelers are willing to pay extra to use MHs when multiple modes are easily accessible through a well-designed physical location supported by a digital platform. Existing studies predominantly focus only on their role as physical transfer points (i.e. transit hubs), where the main objective is to optimally choose the hub locations that enhance accessibility (Petrović et al.,, 2019; Frank et al.,, 2021; Aydin et al.,, 2022). However, these studies often overlook the potential benefits of MHs on operator strategies. For example, MHs provide a location for stationing on-demand vehicles, fueling or charging electric vehicles, and provide a means for two operators to have controlled cost transfers (i.e. only subsidize trips that start or end at MHs). For example, an integration between Long Island Railroad (LIRR) and Uber may be costly if LIRR subsidized all Uber trips to all stations, but setting boundaries for specific stations through bundled tickets purchased for Uber and LIRR can produce partnerships in a more cost-effective manner that can suit both parties.
Effective service integration requires joint operational planning among operators to improve traveler experiences. Recent studies, such as Xanthopoulos et al., (2024), partially address this gap by considering passenger preferences in determining optimal locations and capacities of mobility hubs. Nevertheless, these works still do not fully explore how such hub characteristics as location, capacity, pricing structures, and subsidies can directly shape operators’ key decisions, including fleet allocation, service area planning, and profitability. In reality, MHs can serve as powerful control levers within public cyberphysical platforms, critically influencing both demand-side traveler choices and supply-side operator strategies. This dual influence remains largely underexplored in the current literature.
Recent advancements in MaaS modeling provide a feasible way in addressing the complex interactions between travelers and operators. Bilevel frameworks and many-to-many stable matching games (Liu and Chow,, 2024; Yao and Zhang,, 2024; Liu et al.,, 2024) have effectively captured market dynamics between travelers and operators. However, these models generally do not explicitly incorporate mobility hubs as integral network gateways. This gap highlights the need to systematically integrate mobility hub considerations into MaaS modeling frameworks, particularly in contexts involving complex interactions between public and private operators, such as providing first-mile and last-mile services to commuter rail stations. In addition, high computational complexity associated with the MaaS models prevents larger scale deployment, which is often associated with MH design scenarios.
We fill this gap in the literature by proposing a mathematical model that optimizes the location and capacity of shared multimodal mobility hubs while maximizing travel utility in urban areas, accounting for multimodal trips and the cost transfers or subsidies between operators. In this context, operators participating in the MHs are part of a MaaS platform, which can be operated by a lead fixed route transit operator. To address computational complexity issues and scalability limitations inherent in existing multimodal flow assignment methods, we introduce the Perturbed Utility Route Choice (PURC)-based assignment game, based on the route choice framework from Fosgerau et al., (2022). Unlike path-enumeration approaches, the PURC framework uses link-level utilities to model multimodal route choices efficiently. This approach facilitates scalable and realistic MaaS market analyses involving mobility hubs. A case study for the neighborhoods along three LIRR stations in Suffolk county, NY, illustrates the application of the proposed mobility hub design model on a real-world scale. In addition, the case study provides valuable insights in how a MH can act as leverage in controlling the amount of subsidies in improving transit station usage.
The paper is structured as follows. We present a literature review of the studies focusing on MH and multi-modal mobility network design games in the following section. We then explain the details of the methodology. Afterwards, we present the LIRR based case study to illustrate the model application and discuss the nuance of the service strategies involved in establishing mobility hubs. Finally, we conclude the study and discuss how the work can be extended.
2 Literature Review
Multimodal systems and mobility hub related studies cover several knowledge fields including network pricing modeling, facility location modeling, discrete choice modeling, spatial analysis, social equity evaluations, game-theory based stable matching, and many more. We do not intend to provide an exhaustive list of literature that covers all aspects of the related studies. Instead, we mainly focus on methodological works that explore the interactions among different participants when designing cooperative multimodal services (i.e. MaaS) or MH systems.
2.1 MaaS platform models
MaaS platforms integrate various transportation modes into bundles that enable more seamless trips involving multiple service operators. A MaaS platform allows users to purchase trips consisting of combinations of mobility options in a single transaction. The pricing and service schemes have attracted considerable attention in recent years (van den Berg et al.,, 2022). Bertsimas et al., (2020) proposed a framework that jointly optimizes schedule frequencies and service pricing in a multi-modal transit network. Given a set of multimodal trip paths, service providers optimize their service frequencies to minimize system-wide wait time under budget constraints, while passengers select their preferred routes based on a nested logit model.
Alternatively, two-sided market matching mechanism is adopted to model MaaS service design problems. Djavadian and Chow, (2017) pioneered the two-sided market mechanism into MaaS modeling. Rasulkhani and Chow, (2019) and Pantelidis et al., (2020) expanded the stable matching approach with transferable utilities to model the interaction of fixed-line service providers and travelers as a two-sided market, first as a many-to-one assignment game and then as a many-to-many assignment game. Travelers select their desired paths, which may involve multiple operators, while service operators determine whether to operate a service link and set service prices. Liu and Chow, (2024) further extended this framework to incorporate mobility-on-demand (MOD) services with congested links for accessing MOD. An exact solution algorithm was proposed to identify stable outcomes and determine the level of subsidy required from the platform to the operators to stabilize empty cores, highlighting the additional resources needed to operate MaaS platforms. Building on the many-to-many matching framework, Yao and Zhang, (2024) proposed a different design in which the MaaS platform acts as an intermediary which purchases capacity from service providers and offers service bundles to travelers with OD-based pricing. Additionally, non-MaaS travelers are considered within the congested network.
Bilevel structures are also frequently used to formulate MaaS platform design problems. The upper level typically represents decisions made by the service platform or providers, while the lower level captures the decisions of other participants, such as travelers. Xi et al., 2024a proposed a single-leader multi-follower game (SLMFG) in which the MaaS platform is the leader, and travelers and service providers are followers. The platform determines service bundles and pricing, while travelers and service providers choose their participation to optimize their respective objectives. Building on this work, Xi et al., 2024b expanded the framework into a multi-leader multi-follower game (MLMFG) that considers competition among platforms in an electric MaaS (E-MaaS) ecosystem. Huang et al., (2024) used an SLMFG to model the interaction between a regulator and multiple service providers: the regulator provides path-based subsidies at the upper level, while non-cooperative service providers set link prices to maximize their own profits at the lower level. Pinto et al., (2020) adopted a bilevel structure to address a resource allocation problem, designing a multimodal system with autonomous vehicle (AV) fleets while capturing time-dependent mode choice behavior. Similarly, Bandiera et al., (2024) formulated a bilevel problem in which service operators determine optimal strategies at the upper level and traveler behaviors are modeled at the lower level. Instead of a centralized operator, each service provider maximizes its own profit, resulting in a Nash equilibrium at the upper level.
The deterministic assignment game in (Liu and Chow,, 2024; Yao and Zhang,, 2024) was extended to a probabilistic approach by Liu et al., (2024) to reflect stochastic coalition choice. This choice is not identical to route choice. The assignment model adopts a similar path-based framework as the capacitated SUE model from Bell, (1995), but reflects the choices of the travelers and operators on which coalitions to form. This difference from a conventional route choice model is reflected in the added operator utility term in the lower-level objective function. Because the conventional SUE model requires path-based solutions, path set generation is a prerequisite of the solution process for the stochastic assignment game in Liu et al., (2024) as well. However, high quality path set generation is a non-trival task as discussed in (Prato,, 2009). Furthermore, the logit-based SUE in the lower level of a bilevel optimization makes it hard to integrate into a single level problem for solving to global optimality. As such, the path-based stochastic assignment game is not well-suited for large-scale applications.
2.2 Mobility Hubs
Mobility hubs can serve as key gateways that facilitate multi-modal mobility services. However, studies related to mobility hubs mainly focus on facility location and resource allocation, i.e. aspects of traditional transit hub design in which interactions between operators are ignored. Previous research has primarily focused on locating mobility hubs to improve accessibility and reduce social inequity, with the involved facilities typically centered around a single mode. Caggiani et al., 2020a and Caggiani et al., 2020b proposed models to optimally place bike-share stations to enhance spatial fairness in mobility access. Similarly, Duran-Rodas et al., (2021) developed a model to site bike-share stations with the goal of minimizing spatial inequity. Other studies on hub location decisions also treat social welfare as a central research focus (Banerjee et al.,, 2020; Aydin et al.,, 2022). Frank et al., (2021) improve rural accessibility by using mobility hubs as gateways for multimodal trip itineraries.
Although numerous models have been proposed for designing mobility hubs, only a few studies consider the interactions among mobility service participants. Nair and Miller-Hooks, (2014) was a pioneer in developing hub location models that incorporate traveler mode choices. They proposed a bilevel framework, with the hub operator as the upper-level decision maker and travelers as the lower-level decision makers. Ma et al., (2019) provide a ridesharing operation strategy that prioritize interconnections with transit networks. Though not specifically designed for mobility hub application, the case study uses LIRR stations as key points for deploying proposed ridesharing dispatch and fleet repositioning strategy, showing the potential of using multimodal services for last-mile trips. Xanthopoulos et al., (2024) also demonstrate the importance of considering user preferences in determining optimal hub locations and capacities. Their framework decomposes hub decisions and traveler choices into multiple modules, and a customized metaheuristic was developed to apply the framework at a city scale.
2.3 Research gaps and our contributions
Although numerous studies have developed models for designing multimodal service networks and mobility hubs, explicitly integrating mobility hubs into MaaS optimization frameworks remains a significant and under-explored area. Nair and Miller-Hooks, (2014) and Xanthopoulos et al., (2024) captured the interaction between operator resource allocation decisions and traveler path choices. However, such analyses primarily treat hubs as mere transfer points, without explicitly considering their strategic role in operator decision-making processes. Specifically, service pricing and subsidy between operators are not considered key decision variables, even though they play crucial roles in influencing traveler choices. Xi et al., 2024a ; Xi et al., 2024b proposed frameworks that leverage service capacity and pricing strategies in designing MaaS ecosystems. However, these frameworks do not consider the spatial aspects of resource allocation decisions and traveler itineraries, nor do they incorporate physical infrastructure when deciding service strategies.
The frameworks proposed by Liu and Chow, (2024) and Yao and Zhang, (2024) effectively capture the complex interactions among the MaaS platform, service providers, and travelers while considering flexible service strategies in multimodal networks. Nonetheless, the role of mobility hubs as key levers in MaaS ecosystem design remains unaddressed. Additionally, the above-mentioned studies often require substantial computational resources, hindering their deployments on a city-level scale. These gaps call for further modeling approaches capable of capturing strategic interactions among operators and travelers, enhanced by more scalable computational methods. A path-based stochastic assignment game approach (Liu et al.,, 2024; Liu,, 2024) addresses some of the scalability and capturing of heterogeneous preferences, but faces others with path generation requirements.
This study addresses these gaps by rigorously integrating MH characteristics into MaaS market models, thereby contributing to both theoretical advancement and practical implementation strategies. The main contributions of this study are summarized as follows:
  • 1.
    We propose a bilevel mathematical model for the general MaaS platform design problem, given the network structure. The model aims to maximize MaaS platform revenue or social welfare while capturing the two-sided matching between mobility hub costs and traveler utility at the lower level. The upper-level decision variable is the service subsidy between operators through transfers at the MH, while the lower-level decision variables are link flow and mobility service node capacities.
  • 2.
    We adopt a random utility model (RUM) based formulation to capture the choices of the coalitions between traveler and operator choices in the lower level as proposed by Liu, (2024). We extend it to a link-based formulation using the PURC approach proposed by Fosgerau et al., (2022). The approach offers improved scalability while still capturing link flows jointly determined by service operators and travelers, similar to other RUM formulations (e.g., stochastic user equilibrium, SUE).
  • 3.
    We reformulate the bilevel problem into a single-level problem using the Karush–Kuhn–Tucker (KKT) conditions of the lower-level problem. A gap function-based approach is proposed to accelerate the solution process. Because of the simplicity of the PURC-based quadratic program in the lower level, a global optimum can be attained within a reasonable time frame when applied to large-scale cases.
  • 4.
    We apply the model to a real-world case study based on several LIRR stations and their surrounding areas to demonstrate the computational efficiency. Various operational schemes are also tested to demonstrate the impact of decisions such as pricing and subsidies. The case study provides insights for real-world policy-making in establishing mobility hubs to promote multimodal trips.
3 Proposed mobility hub platform design methodology
We first present the multimodal network structure and service assumptions. We then present the model formulations in detail. The model formulation for MH platform design takes the general stochastic MaaS assignment game from (Liu et al.,, 2024) and re-formulates it within an equivalent PURC framework. The goal is to provide a scalable framework that jointly decides the service decisions from the operator side and path choices from the customer side, while the platform sets service subsidies.
3.1 Network structure and model assumption
The general notation for MaaS platform design used for the MH platform design follows a multicommodity flow network design problem. The multimodal network $G$, denoted by $(N,A)$, consists of all nodes $N$ and links $A$, serving all origin-destination (OD) groups $S$. Origins and destinations are aggregated into centroids $N_{S}\subset N$. The services include a single fixed route transit (FT) operator acting as the platform for designing the MHs, and a set of MOD services that can choose to participate as feeders for these MHs. The model can be trivially modified to include fixed route buses run by other operators as feeders as well, but for simplicity in notation we assume this MH system deals only with a single fixed route platform/operator and one or more MOD potential feeder operators. The FT operator $f$ belongs to a singleton set $F$ and provides service on its own subnetwork $G_{f}$, which consists of node set $N_{f}\subseteq N_{F}\subseteq N$ and link set $A_{f}\subseteq A_{F}\subseteq A$. Similarly, each MOD fleet $m$ belongs to the MOD operator set $M$, with each operating a subnetwork $G_{m}$ consisting of node set $N_{m}\subseteq N_{M}\subseteq N$ and service link set $A_{m}\subseteq A_{M}\subseteq A$. A MOD service link represents a pickup and origin and drop-off at destination, where the link attributes reflect average performance of the fleet serving this request. Congestion effects, such as waiting for service, are captured by a node capacity at the origin. These ”in-platform” links are store-and-forward links with service queues. Meanwhile, each MOD operator $m$ decides on the node capacities $z_{i}$ for all nodes $i$ belonging to node set $N_{m}$.
A dummy subnetwork $G_{O}$ is used to represent travel outside of the MH platform’s multimodal services (e.g., private car, telecommuting, or an alternative platform), i.e. ”out-of-platform” links. A set of transfer links $A_{0}$ are used to connect centroids $N_{S}$ with service nodes $N_{F}\cup N_{M}\cup N_{O}$, as well as connections between layers of subnetworks to construct the overall multimodal network. Specifically, links connecting other layers to a MOD subnetwork layer are called MOD access links. Fig. 1 illustrates the multimodal network structure. Travelers choose their preferred paths encompassing combinations of services across the network for each OD pair $s\in S$, which may use the platform to access the hub and taken the fixed route transit to the destination (and vice versa), use their own modes to get to the transit station to go to the destination, or use their own mode entirely to get to the destination.
In the case of a mobility hub-oriented system, we consider a platform operator, which may or may not be a leading fixed-route transit operator, working with MOD operators to provide services at designated MHs through a subset of nodes $H\subset N$. These operators may be micromobility providers, ridehail, or microtransit services. The hubs in this setting represent cyberphysical gateways in which participating MOD operators that pick up or drop off passengers purchasing trip bundles within these geofenced hubs could potentially be subsidized. Potential MOD feeder services are introduced that are centered around each hub $h\in H$. Direct access between hubs and surrounding service nodes is represented by a set of added links $A_{H}$. The total flow of all direct links $A_{h}$ leading to hub $h$ shall not exceed the capacity of the hub.
Figure 1: Network illustration
As shown in Fig. 1, we separate subnetworks to be ”in-platform” and ”out-of-platform”. We focus on the service strategy decisions of operators in-platform, including link-based prices and service capacities. Since out-of-platform includes all other options out there (including driving, walking, telecommuting), they are assumed to be uncapacitated and we do not explicitly model competition between the designed platform and other external modes. Instead, we focus on the cooperative game between operators in-platform.
Inside the platform, the FT operator acts as the regulator of a MH platform that engages with MOD services to connect travelers to the FT subnetwork. Therefore, only the price of the MOD service links in the in-platform subnetworks are part of the decision variables. Other decision variables include the service capacities at either the MH or the MOD service nodes within the in-platform subnetworks. Table 1 provides the full list of notations. In addition, we list the model assumptions as follows. To avoid confusion, we define travel cost as the non-monetary penalties associated with links such as travel time. Service price is the monetary values required for using each link. Both terms are converted to utilities.
Table 1: Model Variables and Parameters
Notation Description
Sets
NN Set of all nodes in the network
AA Set of all links in the network
$S$ Set of origin-destination (OD) pairs and population segments
$N_{S}\subset N$ Set of centroid nodes (OD origins/destinations)
$F$ Set of fixed route transit (FT) operators
$N_{F}\subseteq N$ Node set of all FT operators
$A_{F}\subseteq A$ Link set of all FT operators
$N_{f}\subseteq N_{F}$ Node set for FT operator $f\in F$
$A_{f}\subseteq A_{F}$ Link set for FT operator $f\in F$
$M$ Set of MOD (Mobility on Demand) operators
$N_{M}\subseteq N$ Node set of all MOD operators
$A_{M}\subseteq A$ Link set of all MOD operators
$N_{m}\subseteq N_{M}$ Node set for MOD operator $m\in M$
$A_{m}\subseteq A_{M}$ Link set for MOD operator $m\in M$
$A_{O}$ Link set for out-of-platform travel
$A_{0}$ Transfer links between centroids and service nodes, and between layers
$H\subset N$ Set of tentative mobility hubs
$A_{H}$ Added links connecting hubs and service nodes
$A_{h}$ Direct links leading to hub $h\in H$
$A^{-}_{i}$ Set of outbound links of node $i$
$A^{+}_{i}$ Set of inbound links of node $i$
Input Variables (Parameters)
$d_{l}$ Length of link $l\in A$
$c^{t}_{l}$ Traveler cost per unit of link $l\in A$
$c^{o}_{l}$ Operator cost per unit of link $l\in A$
$q_{s}$ OD demand for each OD pair $s\in S$
$z_{i}$ Capacity of MOD node $i\in N_{M}$
$c_{i}$ Per capacity cost of MOD node $i\in N_{M}$
$a_{i,l}$ Node-link incidence: 1-1 if origin, +1+1 if destination, 0 otherwise
$\hat{p}_{l}$ Cap of service price for link $l\in A$
$\bar{q}$ Average OD demand
$\alpha$ Weight of operator utility
Decision Variables
$x_{s,l}$ Normalized flow on link $l\in A$ for OD pair $s\in S$
$v_{i}$ Proportion of maximum capacity opened for MOD node $i\in N_{M}$
$p_{l}$ Service price on link $l\in A$; if lower than operator’s price, reflects a subsidy from the platform
List of assumptions:
  • 1.
    The proposed model only considers strategic planning perspective. Therefore, the model only involves static flow assignment.
  • 2.
    The MH platforms are centrally controlled, which involves the subsidy decisions.
  • 3.
    All cost-related terms are link-additive.
  • 4.
    Utilities are transferable between travelers, operators, and the platform.
  • 5.
    Travel cost of the FT service links and out-of-platform links are fixed.
  • 6.
    Transfer links between centroids and FT nodes have fixed travel cost including wait and walk time, which represents the overall cost of accessing FT services.
  • 7.
    MOD access links are store-and-forward links with service queues, meaning that the access flow remains uncongested until capacity is reached. Once capacity is exceeded, excess flows are diverted to other links with available capacity.
  • 8.
    All FT service links and links in the out-of-service subnetwork have fixed travel cost and are uncapacitated. Congestion from background traffic are directly incorporated into the travel costs. If an out-of-service link reflects multiple modes, its travel cost reflects the expected minimized disutility option of those modes. Alternatively each parallel mode may be modeled separately.
  • 9.
    Operating cost of MOD services consists of service link costs and access capacity allocation costs.
3.2 Lower-level traveler-operator coalition choice model using PURC based structure
The MH platform design problem is adapted from the stochastic MaaS assignment game model in Liu et al., (2024). The assignment game is decomposed into a bilevel problem where the upper level involves platform decisions, such as pricing and subsidies, while the lower level involves the joint decisions of the travelers and mobility operators. In the adaptation to the MH platform design problem, the platform is also the FT operator making decisions about how much to subsidize the MOD operators in the upper level, while the lower level determines the MH and MOD capacities and traveler flows.
Instead of using the path-based SUE-style formulation in Liu et al., (2024), we adapt the link-based PURC model proposed by Fosgerau et al., (2022) to capture the probabilistic matching mechanism. Eq. (1) is the link-based stochastic assignment problem $L_{1}$ based on the PURC concept for a MH platform.
$L_{1}\min_{\{x_{l,s},\,v_{i},\}}\quad$ $\Phi_{1}=\sum_{l\in A}\sum_{s\in S}\bar{q}d_{l}x_{l,s}^{2}+\sum_{l\in A}\sum_{s\in S}\bar{q}d_{l}\left(p_{l}+c^{t}_{l}\right)x_{l,s}+\alpha\left(\sum_{l\in A}\sum_{s\in S}d_{l}c^{o}_{l}x_{l,s}q_{s}+\sum_{i\in N_{M}}z_{i}c_{i}v_{i}\right)$ (1a)
s.t. $\sum_{l\in A}a_{i,l}x_{l,s}=\begin{cases}-1,\quad\forall i=o,\\ 1,\quad\forall i=d,\\ 0,\quad\forall i\in N\setminus\{o,d\}\end{cases}\quad\forall s=(o,d)\in S,\,\forall i\in N$ (1b)
$\sum_{l\in A^{+}_{0,i}}\sum_{s\in S}x_{l,s}q_{s}\leq z_{i}v_{i}\quad\forall i\in N_{M}$ (1c)
$\sum_{l\in A^{+}_{i,h}}\sum_{s\in S}x_{l,s}q_{s}\leq z_{i}v_{i}\quad\forall h\in N_{H}$ (1d)
$0\leq x_{l,s}\leq 1\quad\forall l\in A,\,\forall s\in S$ (1e)
$0\leq v_{i}\leq 1\quad\forall i\in N_{M}$ (1f)
The assignment model $\Phi_{1}$ serves as a lower level problem to the corresponding assignment game, but can also be run as an independent model for determining a multimodal network design problem with stochastic assignment, such as determining the welfare impact of an existing pricing design of a traditional transit hub without MH subsidy.
The decision variables are the link flow assignment $x_{l,s}$ and MOD node capacity $v_{i}$. The service price $p_{l}$ is the decision variable in the upper level model, which is presented in the following subsection. In $\Phi_{1}$, $p_{l}$ is treated as an input. Eq. (1b) ensures the link flow conservation. For each OD pair $s=(o,d)$ with $o$ being the origin and $d$ being the destination, the source node $o$ initiates all normalized flow, and the sink node $d$ terminates all normalized flow. Eq. (1c) controls the capacity and congestion effect of the platform service node. All access links flowing into the service nodes are bounded by the assigned capacity the operator determines. Eq. (1d) is dedicated to mobility hub facilities. For all service flows ending at mobility hubs, the sum of flow needs to be capped by the assigned mobility hub capacity. Eqs. (1e - 1f) are the bounds for the two decision variables.
The objective consists of three terms: the perturbed utility, traveler utility, and the operator utility. Introduced by Fosgerau et al., (2022), the PURC model follows the general RUM structure with link-based property to capture the dispersion of flows across a network. The flow assignment variable $x_{l,s}$ represents the probability of choosing link $l$ for travelers in OD pair $s$. The first term in Eq. (1a) depicts the dispersion effect. The second term is the sum of traveler utility, which depicts the sum of all link utilities. We simply the term by only involving a general cost element and a service price element. It can be expanded to involve all utility related elements such as externalities, as long as they can be converted to utility values. For a comprehensive explanation of the disperse effect and the utility features that can be included, readers can refer to Fosgerau et al., (2022) for more information. The original PURC objective is constructed by combining the first and second terms. We show this in Proposition 1.
Proposition 1.
The first and second terms combined in Eq. (1a) are equivalent to the PURC objective proposed by Fosgerau et al., (2022).
Proof.
The first and second term of Eq. (1a) can be written as Eq. (2).
$\min_{\{x_{l,s},\,v_{i},\}}\quad\Phi_{1}=\bar{q}\sum_{s\in S}d_{l}(x_{l,s}^{2}+\left(p_{l}+c^{t}_{l}\right)x_{l,s})$ (2)
The coefficient $\bar{q}$ is a constant. The quadratic term is the perturbed term $F(x)$, and the linear term is the link-additive utility term $U(x)$. Both terms are weighted by traversed link length $d_{l}$. The structure is therefore identical to the PURC objective proposed in Fosgerau et al., (2022) by multiplying with a constant $\bar{q}$. ∎
We introduce the third term to form the coalition between travelers and operators in a similar fashion to the model proposed in chapter 4 of Liu, (2024). In Proposition 2, we show that $\Phi_{1}$ provides a flow assignment model with a traveler-operator coalition under the rule of PURC.
Proposition 2.
$\Phi_{1}$ yields the link flow assignment under PURC framework with traveler and operator coalition choice.
Proof.
We divide Eq. (1a) by the constant term $\sum{q_{s}}/{|S|}$ and write it as Eq. (3).
$\min_{\{x_{l,s},\,v_{i},\}}\quad\Phi^{\prime}_{1}=\sum_{s\in S}\sum_{l\in A}d_{l}x_{l,s}^{2}+\sum_{s\in S}\sum_{l\in A}d_{l}\left(p_{l}+c^{t}_{l}+\alpha c^{o}_{l}\frac{q_{s}}{\bar{q}}\right)x_{l,s}+\alpha\left(\sum_{i\in N_{M}}\frac{z_{i}c_{i}v_{i}}{\bar{q}}\right)$ (3)
The Lagrangian of $\Phi^{\prime}_{1}$ is written as Eq. (4).
$\min_{\{x_{l,s},\,v_{i},\}}\quad L^{\prime}_{1}$ $=\sum_{s\in S}\sum_{l\in A}d_{l}x_{l,s}^{2}+\sum_{s\in S}\sum_{l\in A}d_{l}\left(p_{l}+c^{t}_{l}+\alpha c^{o}_{l}\frac{q_{s}}{\bar{q}}\right)x_{l,s}+\alpha\left(\sum_{i\in N_{M}}\frac{z_{i}c_{i}v_{i}}{\bar{q}}\right)$ (4a)
$+\sum_{s\in S}\sum_{i\in N}\mu_{i,s}(\sum_{l}a_{i,l}x_{l,s}-f_{i,s})$ (4b)
$+\sum_{i\in N_{M}}\lambda_{i}(\sum_{l\in A^{-}_{0,i}}\sum_{s\in S}x_{l,s}q_{s}-z_{i}v_{i})$ (4c)
$+\sum_{l\in A}\sum_{s\in S}\beta_{l,s}(x_{l,s}-1)+\sum_{i\in N_{M}}\pi_{i}(v_{i}-1)$ (4d)
The optimal solution of $x_{l,s}$ is obtained by having $\frac{\partial L^{\prime}_{1}}{x_{l,s}}=0$ for all chosen links $l$ for OD pair $s$.
$\frac{\partial L^{\prime}_{1}}{x_{l,s}}$ $=2d_{l}x_{l,s}+d_{l}\left(p_{l}+c^{t}_{l}+\alpha\frac{c^{o}_{l}q_{s}}{\bar{q}}\right)$ (5a)
$+\sum_{i\in{l^{+},l^{-}}}\mu_{i,s}a_{i,l}+\sum_{i\in{l^{+},l^{-}}}\lambda_{i}q_{s}+\beta_{l,s}=0$ (5b)
Because the chosen links will have strictly positive flows, the $v_{i}$ connecting those links also fulfill the optimality conditions.
$\frac{\partial L^{\prime}_{1}}{v_{i}}$ $=\alpha\frac{z_{i}c_{i}}{\bar{q}}-\lambda_{i}z_{i}=0$ (6a)
By re-writing Eq. (6), we obtain $\lambda_{i}=\frac{\alpha c_{i}}{\bar{q}}$. By plugging it in Eq. (5), we have Eq. (7).
$\frac{\partial L^{\prime}_{1}}{x_{l,s}}$ $=2d_{l}x_{l,s}+d_{l}\left(p_{l}+c^{t}_{l}+\alpha\frac{c^{o}_{l}q_{s}}{\bar{q}}\right)$ (7a)
$+\sum_{i\in{l^{+},l^{-}}}\mu_{i,s}a_{i,l}+\sum_{i\in{l^{+},l^{-}}}\alpha c_{i}\frac{q_{s}}{\bar{q}}+\beta_{l,s}=0$ (7b)
Eq. (7) can be rewritten as Eq. (8).
$x_{l,s}=-\frac{1}{2}\left(p_{l}+c^{t}_{l}+\alpha c^{o}_{l}q_{s}\frac{|S|}{\sum{q_{s}}}\right)-\frac{1}{2d}(\sum_{i\in{l^{+},l^{-}}}\mu_{i,s}a_{i,l}+\sum_{i\in{l^{+},l^{-}}}\alpha c_{i}q_{s}\frac{|S|}{\sum{q_{s}}}+\beta_{l,s})$ (8)
Eq. (8) depicts the flow assignment decisions for all links $l$ with positive flows for each OD pair $s$. The first term consists of the link disutility from both travelers and operators. The link disutility from the operator side is weighted by $\alpha$ and $q_{s}/\bar{q}$. $\alpha$ is a general coefficient that provides a different weight towards operator disutility compared to travelers. A higher weight leads to higher impact on operator perspective. $q_{s}/\bar{q}$ is an additional weight that considers the impact from different OD pair volumes. The flow assignment variable $x_{l,s}$ is considered as the possibility of choosing link $l$ for OD group $s$. Each OD pair shall be treated equally from a traveler perspective to reflect stochastic user-equilibrium (SUE) behavior from the perspective of the coalitions of travelers and operators. The weight $q_{s}/\bar{q}$ is the lever that place more focus on OD pairs with higher OD volume. For OD pair with higher than average flow, this causes the OD pair $s$ having more impact on final assignment decision for operators. If there are two OD pairs $(0,1)$ and $(0,2)$. $(0,1)$ has 1 unit flow and $(0,2)$ has 2 units. Both ODs would flow through a mobility hub. On the traveler side, each OD is equally treated for flow assignment. However, when it comes to operators, OD pair $(0,2)$ has a higher weight in the resource allocation decision from a system-wise perspective. Therefore, a weight of 0.67 is placed for node cost disutility when considering OD pair $(0,1)$, while the weight towards OD pair $(0,2)$ becomes 1.33.
The second term consists of node specific values except the capacity Lagrangian $\beta_{l,s}$. The first element represents the magnitude of attraction of link $l$ in OD pair $s$, which is measured by the difference between the source node Lagrangian and the sink node Lagrangian. The second element represents the disutility from the node cost for MOD operators. For links that do not belong to the access link set leading to MOD nodes $N_{M}$, the second element is dropped. For the the access links leading to MOD nodes, the node cost disutility from the operator side is counted towards the flow assignment, which is also weighted by $q_{s}/\bar{q}$.
Eq. (8) mimics the structure of the optimal link-flow decision variable with non-zero value shown in Fosgerau et al., (2022). This concludes the proof. ∎
Bell, (1995) proved that the congestion effect can be reflected by the capacity Lagrangian in a capacitated network, which is also shown in Liu et al., (2024). We follow the same logic to show that the congestion effect can also be evaluated by capacity Lagrangians in the PURC based assignment model.
Proposition 3.
The congestion effect of waiting for service at a capacitated MOD node $i\in N_{M}$ is captured by the corresponding Lagrangian $\lambda_{i}$.
Proof.
Assume a node $i$ is at capacity ($v_{i}=1$). If there is an additional flow $\delta_{s}$ for OD pair $s$ on link $l$, with $i$ being the sink node, the capacity constraint Eq. (1c) is therefore exceeded by $\frac{\delta_{s}q_{s}}{z_{i}}$. Assume the additional flow switches to link $l^{\prime}$ while the chosen downstream links have identical costs compared to the chosen downstream links sourcing from node $i$. The change of objective value measured by Eq. (1a) is written as Eq. (9).
$d\Phi_{1}=\bar{q}(d_{l}(2x_{l,s}+\delta_{s})-d_{l^{\prime}}(2x_{l^{\prime},s}+\delta_{s}))+\bar{q}\delta_{s}(d_{l}p_{l}-d_{l^{\prime}}p_{l^{\prime}}+d_{l}c^{t}_{l}-d_{l^{\prime}}c_{l^{\prime}}^{t}))+\alpha\delta_{s}(d_{l}c^{o}_{l}-d_{l^{\prime}}c^{o}_{l^{\prime}})q_{s}$ (9)
As shown in Eq. (10), when the right-hand-side of the node capacity constraint Eq. (1f) increases by 1, the objective value of the KKT in Eq. (4) increases by $\pi_{i}$. Therefore, we can rewrite Eq. (9) to Eq. (11).
$\frac{d\Phi_{1}}{\delta_{s}q_{s}/z_{i}}=\frac{z_{i}\bar{q}}{q_{s}}(d_{l}(2x_{l,s}+\delta_{s})-d_{l^{\prime}}(2x_{l^{\prime},s}+\delta_{s}))+\frac{z_{i}\bar{q}}{q_{s}}(d_{l}p_{l}-d_{l^{\prime}}p_{l^{\prime}}+d_{l}c^{t}_{l}-d_{l^{\prime}}c_{l^{\prime}}^{t}))+\alpha z_{i}(d_{l}c^{o}_{l}-d_{l^{\prime}}c^{o}_{l^{\prime}})=\pi_{i}$ (10)
$\frac{\bar{q}}{q_{s}}(d_{l}(2x_{l,s}+\delta_{s})-d_{l^{\prime}}(2x_{l^{\prime},s}+\delta_{s}))+\frac{\bar{q}}{q_{s}}(d_{l}p_{l}-d_{l^{\prime}}p_{l^{\prime}}+d_{l}c^{t}_{l}-d_{l^{\prime}}c_{l^{\prime}}^{t}))+\alpha(d_{l}c^{o}_{l}-d_{l^{\prime}}c^{o}_{l^{\prime}})=\frac{\pi_{i}}{z_{i}}$ (11)
If link $l$ and $l^{\prime}$ both having 1 unit of length, Eq. (11) is equivalent to Eq. (12).
$x_{l,s}-x_{l^{\prime},s}=-(p_{l}-p_{l^{\prime}})-(c^{t}_{l}+\frac{\pi_{i}q_{s}}{z_{i}\bar{q}}-c_{l^{\prime}}^{t})-\alpha(c^{o}_{l}-c^{o}_{l^{\prime}})\frac{q_{s}}{\bar{q}}$ (12)
Eq. (12) shows that the congestion effect caused by the capacitated node can be evaluated by the Lagrangian of the node capacity constraint. When the switched links are identical in length, the congestion effect equals to $\frac{\pi_{i}q_{s}}{z_{i}\bar{q}}$. This concludes the proof. ∎
The lower-level assignment model is a convex quadratic programming (QP) problem, which can be easily solved with off-the-shelf solvers at scale. Moreover, it exhibits characteristics that allow us to solve the bilevel problem to global optimality as well.
3.3 Upper level model formulation
We present the upper-level problem in this subsection. The decision-maker of the upper-level problem is the MH platform, which determines the optimal subsidy strategy for the mobility services. The bilevel structure follows the Stackelberg equilibrium principle, with the MaaS platform as the leader that sets centralized pricing strategies, while travelers and operators are followers who make decisions to minimize their disutilities. We use revenue maximization as the upper-level objective. However, other equity or social welfare focuses objective can be applied for the upper level objective. From a revenue maximization perspective, the mobility hub design problem can be formulated as Eq. (13).
$\max_{\mathbf{p}}\quad$ $\Phi_{0}=\sum_{s}\sum_{l}d_{l}p_{l}\,x_{s,l}\,q_{s}$ (13a)
s.t. $\mathbf{x}=\arg\min_{\mathbf{x}}\;\Phi_{1}(\mathbf{p})$ (13b)
$0<p_{l}\leq\hat{p},\quad\forall l\in A_{M}$ (13c)
The upper-level problem finds the optimal pricing strategy that maximizes the total revenue from the MH platform users while the service flows and capacities are determined by the lower-level problem. The model assumes that $\hat{p}$ reflects the existing price offered by the MOD operator, and $p_{l}$ is the subsidized price at the link level. Other pricing schemes can be considered as well, such as keeping the same subsidy rate for all trips arriving at a particular MH (by indexing price by MH), or having the same subsidy rate per operator $m$ by indexing price by operator instead.
We can also change the objective to reflect maximization of passenger-miles with an explicit term for subsidies. Assuming there is a fixed subsidy package $R$ provided to the platform operator and the service price is fixed as $p^{{}^{\prime}}$, then the upper-level problem would becomes a flow maximization objective that optimally allocate subsidies. Therefore, Eq. (13) can be rewritten as Eq. (14).
$\max_{\mathbf{r}}\quad$ $\Phi_{0}=\sum_{s}\sum_{l}d_{l}\,x_{s,l}\,q_{s}$ (14a)
s.t. $\sum_{l}d_{l}r_{l}x_{l,s}q_{s}=R$ (14b)
$\mathbf{x}=\arg\min_{\mathbf{x}}\;\Phi_{1}(\mathbf{p^{{}^{\prime}}-r})$ (14c)
$r_{l}\geq 0,\quad\forall l\in A_{M}$ (14d)
The objective of Eq. (14) is flow maximizaation. The service price travelers need to pay in the lower-level is therefore $p^{{}^{\prime}}-r$ as shown in Eq. (14c). In this variant, the total amount of subsidy is fixed. It reflects a scenario where the policymaker seeks to evaluate trade-offs between level of subsidy and flow, particularly relevant for public MH platforms.
Aside from the PURC modifications, the bilevel problem presented in these two subsections feature several nontrivial differences from the generic MaaS assignment game model from (Liu and Chow,, 2024) and (Liu et al.,, 2024). These include:
  • 1.
    In-platform links only include the ones inbound and outbound from MHs; the consequence is the capacity designs focus only on which last-mile service regions the MOD operators would serve (capacities approaching zero suggests not serving those regions) and the capacities for the MHs (capacities approaching zero here suggests closing the MH).
  • 2.
    There are out-of-platform links not just connecting centroid OD pairs, but also connecting centroids to the FT operators’ stations as out-of-platform access links for travelers choosing to use the FT service without engaging with the platform.
  • 3.
    Whereas the stochastic assignment game in (Liu et al.,, 2024) lacks a subsidy and the deterministic one in (Liu and Chow,, 2024) only subsidizes a system that has no stable outcome otherwise, the proposed model uses subsidy as the primary upper level pricing decisions, assuming that the FT operator and MOD operators already have set pricing. As shown in (Liu et al.,, 2024), the stochastic assignment game is guaranteed to have a stable outcome with a single cost allocation corresponding to the optimal flow because of the endogeneity in the coalition choices. A solution in which optimal subsidies are zero suggests there is no need to setup a MaaS platform with fare bundling as there should be sufficient incentive for travelers and operators to interact for last mile access out-of-platform.
  • 4.
    The FT operator serves the dual role of being an operator (with lower level MH capacity decisions) and also a platform (with control of the upper level subsidy decisions).
  • 5.
    The general MaaS platform assignment game assumes there is no preexisting network, so as costs for operators go up, the optimal solution will have everybody move to out-of-platform alternatives. In the proposed model, there is a preexisting FT service, so if costs of operating MHs are too high, out-of-platform flows imply travelers either going to their destinations on their own, or going to the transit stations on their own or via an unsubsidized access mode.
3.4 Solution method
The bilevel programming problem is nonconvex due to the dependency between leader and follower. The bilevel problem can be reformulated into a single-level constrained optimization problem by incorporating the KKT conditions of Eq. (1) as additional constraints into Eq. (13) or Eq. (14). We use Eq. (13) as illustration and the reformulated single-level problem is written as Eq. (15).
$\max_{\mathbf{p,x,v,\mu,\lambda,\beta,\pi}}\quad$ $\Phi^{\prime}_{0}=\sum_{s}\sum_{l}d_{l}p_{l}\,x_{s,l}\,q_{s}$ (15a)
s.t. $\sum_{l}d_{l}r_{l}x_{l,s}q_{s}=R$ (15b)
$0<p_{l}\leq\hat{p},\quad\forall l\in A_{M}$ (15c)
$x_{l,s}(2d_{l}x_{l,s}+d_{l}\left(p_{l}+c^{t}_{l}+\alpha\frac{c^{o}_{l}q_{s}}{\bar{q}}\right)$
$+\sum_{i\in{l^{+},l^{-}}}\mu_{i,s}a_{i,l}+\sum_{i\in{l^{+},l^{-}}}\lambda_{i}q_{s}+\beta_{l,s})=0\quad\forall l\in A,s\in S$ (15d)
$\alpha\frac{z_{i}c_{i}}{\bar{q}}-\lambda_{i}z_{i}+\pi_{i}=0,\quad\forall i\in N_{M}$ (15e)
$\mu_{i,s}(\sum_{l}a_{i,l}x_{l,s}-f_{i,s})=0,\quad\forall i\in N,s\in S$ (15f)
$\lambda_{i}(\sum_{l\in A^{-}_{0,i}}\sum_{s\in S}x_{l,s}q_{s}-z_{i}v_{i})=0,\quad\forall i\in N$ (15g)
$\beta_{l,s}(x_{l,s}-1)=0,\quad\forall l\in A,s\in S$ (15h)
$\pi_{i}(v_{i}-1),\quad\forall i\in N$ (15i)
$\sum_{l\in A}a_{i,l}x_{l,s}=\begin{cases}-1,\quad\forall i=o,\\ 1,\quad\forall i=d,\\ 0,\quad\forall i\in N\setminus\{o,d\}\end{cases}\quad\forall s=(o,d)\in S,\,\forall i\in N$ (15j)
$\sum_{l\in A^{-}_{0,i}}\sum_{s\in S}x_{l,s}q_{s}\leq z_{i}v_{i}\quad\forall i\in N_{M}$ (15k)
$0\leq x_{l,s}\leq 1\quad\forall l\in A,\,\forall s\in S$ (15l)
$0\leq v_{i}\leq 1\quad\forall i\in N_{M}$ (15m)
$\mathbf{p,x,v,\mu,\lambda,\beta,\pi}\geq 0$ (15n)
Eqs. (15d, 15e) are the first order conditions. Eq. (15d) is an augmented first order condition because of the zero-flow links involved in the final solution as indicated in Fosgerau et al., (2022). Eqs. (15f - 15i) are the complementarity slackness conditions. It has been shown (e.g. Dempe, (2002)) that bilevel problems with linear upper level and convex quadratic lower level programs can be reformulated into a single level mathematical program with complementarity constraints (MPCC) that can be solved to global optimality. Sinha et al., (2017) provide an example algorithm based on branch-and-bound.
However, the added KKT multipliers and complementarity constraints pose great scalability challenges, especially when facing a large network. Therefore, it is impractical to solve for a global optimal solution using branch-and-bound. To address this issue, we adopt the gap function-based method proposed by Marcotte and Zhu, (1996) to solve the bilevel problem effectively while still keeping high solution quality. Instead of directly appending the KKT conditions into upper-level constraints, we move part of the high-dimension constraints into the objective function and formulate them as penalty terms. The new penalty added objective and the adjusted constraints are written as Eq. (16).
$\min_{\mathbf{p,x,v,\mu,\lambda,\beta,\pi}}\quad$ $\Phi^{\prime}_{0}=-\sum_{s}\sum_{l}d_{l}p_{l}\,x_{s,l}\,q_{s}+\rho\sum_{l,s}\Lambda_{l,s}$ (16a)
s.t. $\Lambda_{l,s}\geq 0,\quad\forall l\in A,s\in S$ (16b)
Eqs. (15b,15c,15e-15m) (16c)
The penalty term $\Lambda_{l,s}$ is the augmented first order condition Eq. (15d) for each l,s{l,s} pair. The penalty coefficient $\rho$ is a tunable hyperparameter that modifies the magnitude of the penalty. The higher the value of $\rho$, the heavier the penalty when not solved towards the lower-level optima. When $\rho$ is sufficiently high, Eq. (16) is equivalent to minimizing the upper level objective while obtaining the lower level optimum. Therefore, we can also use Eq. (16) to obtain a global optimum of this bilevel problem.
Proposition 4.
There exists a value of $\rho$ for which the solution obtained from solving Eq. (16) converges to the optimal solution obtained from Eq. (15).
Proof.
The equality constraint Eq. (15d) can be written as Eq. (17).
$\Lambda_{l,s}\geq 0,\quad\forall l\in A,s\in S$ (17a)
$\Lambda_{l,s}\leq 0,\quad\forall l\in A,s\in S$ (17b)
We denote the Lagrangian multiplier of Eq. (17b) to be $\eta_{l,s}$. We switch the original objective Eq. (15a) to a minimization problem. We then relax Eq. (17b) to have Eq. (18).
$\min\quad$ $-\sum_{s}\sum_{l}d_{l}p_{l}\,x_{s,l}\,q_{s}+\sum_{l,s}\eta_{l,s}\Lambda_{l,s}$ (18a)
$\Lambda_{l,s}\geq 0,\quad\forall l\in A,s\in S$ (18b)
When $\rho\geq max(\eta)$, the penalty term $\rho\sum\Lambda$ dominate the incentive to violate Eq. (17b). Under this condition, any optimal solution of the original problem is optimal for the penalized problem. The equality constraint Eq. (15d) is recovered at the optimum even though the feasible set only enforces Eq. (18b). This completes the proof. ∎
To further reduce the computational load, we propose another algorithmic improvement. Since the lower-level problem is very efficient to solve due to its convexity, it is easy to obtain the lower-level optimum when p is fixed. We use an iterative update process to exploit this property and reduce the solution time for Eq. (16). The overall proposed algorithm is summarized in Algorithm 1. The hyperparameters $\rho^{0}$, $\psi^{-}$, and $\psi^{+}$ control the rate of penalty update, where $\psi^{+}>1$ and $\psi^{-}<1$. By starting with a small value $\rho^{0}$, Eq. (16) is easier to solve to optimality. This comes with the cost of violating lower-level KKT condition by a large margin ($\sum\Lambda>>\epsilon$). By gradually increasing $\rho$ with $\psi^{+}$, the KKT violation shrinks, leading to the global optimum. However, when $\rho$ becomes too large, the single-level solution process can become computationally intensive. Therefore, we reduce $\rho$ by $\psi^{-}$ to control the step size. Within each iteration, the previously solved $\mathbf{p}$ (or $\mathbf{r}$ if the decision variable is the subsidy) is used to solve the lower-level problem first. This solution, along with all dual values and their bounds, can warm-start the single-level solution algorithm in a tighter space, leading to a shorter runtime. After meeting the convergence criteria or reaching the maximum number of iterations, we output the final values of $(\mathbf{p}^{*},\mathbf{x}^{*},\mathbf{v}^{*})$.
Algorithm 1 Penalty-Based Iterative Solution Method for Mobility Hub Platform Design Problem
1:Input: Initial upper-level variable $\mathbf{p}^{0}$, initial penalty parameter $\rho^{0}$, bound relaxation coefficient $\zeta$, penalty parameter update coefficient $\psi^{-},\psi^{+}$, optimality gap target $\tau$, objective tolerance $\epsilon$, time limit $T$, maximum iterations $K$
[…]
3:repeat
4:  Step 1: For given $\mathbf{p}^{k}$, solve the lower-level problem Eq. (1) to obtain optimal $\mathbf{x}^{k},\mathbf{v}^{k}$
5:  Step 2: Obtain all Lagrangians $\mathbf{\lambda^{k},\mu^{k},\beta^{k},\pi^{k}}$ from the lower level solutions and their current bounds.
6:  Step 3: Load and warm start the solution process for Eq. (16) using $\mathbf{x^{k},v^{k},\lambda^{k},\mu^{k},\beta^{k},\pi^{k}}$ and $\rho^{k}$. Update the bounds for $\mathbf{\lambda^{k},\mu^{k},\beta^{k},\pi^{k}}$ by $\zeta$. Solve until time limit $T$ is reached or the optimality gap is lower than $\tau$. Obtain the new solution $\mathbf{p}$.
7:  Step 4: Check optimality condition:
8:  if Optimality gap $\leq\tau$ then
9:   Go to step 5
10:  else
11:   $\rho^{k}\leftarrow\psi^{-}\rho^{k}$
12:  end if
13:  Step 5: Check penalty condition:
14:  if $\sum\Lambda\leq\epsilon$ or $k\geq K$ then
15:   Break
16:  else
17:   $\rho^{k}\leftarrow\psi^{+}\rho^{k}$
18:  end if
19:  Step 6: $\textbf{p}^{k}=\textbf{p}$
20:  $k\leftarrow k+1$
21:until convergence
22:Solve lower level problem Eq. (1) with $\textbf{p}^{*}$. Obtain $\mathbf{x}^{*},\mathbf{v}$
23:Output: Optimal solution $(\mathbf{p}^{*},\mathbf{x}^{*},\mathbf{v}^{*})$.
4 Numerical experiments
In this section, we provide two sets of experiments to illustrate the formulation and the solution algorithm. We first use a toy network to verify the algorithm and illustrate the model. The second result is tested on a hypothetical multimodal network centered around three LIRR stations, which serve as potential mobility hubs for travelers commuting to New York City from Long Island.
4.1 Illustrative example
A toy network is illustrated in Fig. 2. Three OD pairs are considered: (1, 0), (2, 0), and (3, 0), each with a demand of 100 units over a typical operating period. The MaaS platform is considered for the transit line from station S to node 0, as well as MOD service links provided by a single operator from nodes 1, 2, and 3 to node H to connect to station S. The gray dotted links represents non-MaaS participating services (e.g., privately owned vehicles, walking, other non-MaaS services). For this illustrative case, only the MH node $H$ has the capacity constraint. The MH capacity reflects the supply side resources for supporting a fleet there: dedicated space for temporary stationing, maintenance, fueling, etc.
All input parameters are listed in Table 2. All access and egress links are assigned with a small length value, which is used to prevent infeasible solutions without impacting the flow assignment results. The three MOD service links (A-H, B-H, C-H) are controlled by the MaaS platform to adjust suitable service pricing in the upper level problem, while other services having fixed pricing aggregated with the the travel cost. $c_{l}^{t},c_{l}^{o},p_{l}$ are all measured in monetary values. The lower level problem decides the joint choice of traveler flow and hub capacities.
The toy example is solved on a device with an Intel(R) i7-13705 processor. We directly solve it to global optimality using Gurobi 12.1 and the branch-and-bound algorithm, with the optimality gap being less than $10^{-6}$. All scenarios reaches global optimum under 0.5 seconds.
BENCHMARK SCENARIO: To better understand how the price allocation scheme impact both the MH based service flow and platform revenue, we run a benchmark scenario representing a traditional transit hub where operators may converge for transfers, but no digitalization is setup for cost transfers between the operators. This solution is obtained by simply solving the lower-level model only with a fixed price of \$3\\$3. When MOD service link prices are all fixed to \$3 per unit distance, only link (A, H) is actively used with a flow of 24.47 from node 1. Both nodes 2 and 3 are not served by the MOD service. Total revenue is \$73.4 and the proportion of opened MH capacity is 12.24%.
BASE SCENARIO: The results are summarized in Table 3 under the columns ”Price 1” and ”Flow 1.” OD pair (3, 0) exclusively relies on the direct non-MaaS link, whereas OD pairs (1, 0) and (2, 0) exhibit multimodal routing behavior that combines both direct and MaaS-integrated links. Notably, OD (1, 0) distributes its flow across three routes: the direct link, a non-MaaS path using the transit connection (1S0)(1-S-0), and a MOD-only path via (1AHS0)(1-A-H-S-0), highlighting the benefits to travelers from node 1 when MOD services are introduced. The MaaS-based MOD service links (A, H) and (B, H) are priced at 3.0 and 0.937, respectively, to maximize total revenue. The latter reflects a subsidy of \$2.06 per unit distance for travelers using MOD service from zone 2. For OD pair (3, 0), the direct link in the dummy subnetwork dominates all other options. even when the platform offers free access to the MOD service link (C, H), no travelers would choose to use it nor would the operator serve that area. Due to the overall low usage of MaaS services in this setting, the hub is heavily underutilized, with a capacity set to only 16.9% of the maximum capacity. The total MaaS service revenue in this case is \$90.8, a 23.7% increase over the benchmark. The value of having a cyberphysical MH with enabled subsidy is equal to \$16.6 per operating period.
Remark 1.
Compared to a benchmark transit hub, the subsidy-enabled MH can produce a higher capacity and increased revenues; the value of having a such a MH can be quantified as the difference in the objective value.
ALTERNATIVE SCENARIO: Now consider a scenario where the service price is capped to \$2\\$2 per unit distance. As shown in the ”Price 2” and ”Flow 2” columns of Table 3, more travelers use the MOD service links on link (A, H) due to the lowered cap. For link (B, h), the price level remains the same, although with the price cap the required subsidy is reduced to only \$1.06. Again, no travelers choose link (C, H) due to the dominance of the direct link (3, 0). As a result, the platform sets the capacity to 20.31% of maximum hub capacity to serve MaaS users. The MaaS platform earns a total of \$80. The \$10 dollar loss in revenue due to further constraining the maximum price translates to a higher enrollment in users coming from zone 1 at an overall lower social surplus.
Remark 2.
The model captures the social surplus and ridership impacts of a price cap.
Figure 2: Toy network illustration
Table 2: Network Input and Parameters
Source node Sink node $\mathbf{c_{l}^{t}(\$)}$ $\mathbf{c_{l}^{o}(\$)}$ $\mathbf{\hat{p}_{l}(\$)}$ $\mathbf{d_{l}}$
1 0 7 0 N/A 4
2 0 7 0 N/A 4
3 0 7 0 N/A 4
A H 4 1 3 1
B H 4 1 3 2
C H 4 1 3 3
1 S 6 0 N/A 1
2 S 6 0 N/A 2
3 S 6 0 N/A 3
H S 0 0 N/A 0.1
S 0 6 0 N/A 4
1 A 0 0 N/A 0.1
2 B 0 0 N/A 0.1
3 C 0 0 N/A 0.1
Maximum hub capacity, $\mathbf{z_{i}}$ 200
$\mathbf{\alpha}$ 0.5
Capacity cost, $\mathbf{c_{i}}$ 1
Table 3: Comparison of link prices and flows across base and alternative scenarios, with subsidies in ( )
Link OD Pair $s$ Price 1, $p_{l}$ ($r_{l}$) Flow 1, $x_{l,s}q_{s}$ Price 2, $p_{l}$ ($r_{l}$) Flow 2, $x_{l,s}q_{s}$
(1, 0) (1, 0) N/A 53.07 N/A 49.28
(2, 0) (2, 0) N/A 90.71 N/A 90.71
(3, 0) (3, 0) N/A 100 N/A 100
(1, A) (1, 0) N/A 24.47 N/A 31.30
(A, H) (1, 0) 3 24.47 2 (0) 31.30
(H, S) (1, 0) N/A 24.47 N/A 31.30
(2, B) (2, 0) N/A 9.29 N/A 9.29
(B, H) (2, 0) 0.94 (2.06) 9.29 0.94 (1.06) 9.29
(H, S) (2, 0) N/A 9.29 N/A 9.29
(1, S) (1, 0) N/A 22.46 N/A 19.42
(S, 0) (1, 0) N/A 46.93 N/A 50.72
(S, 0) (2, 0) N/A 9.29 N/A 9.29
Hub Capacity, $v_{i}$ 16.88% 20.31%
Objective value, $\Phi_{0}$ \$90.8\\$90.8 \$80\\$80
4.2 LIRR mobility hubs case study
We further use three LIRR stations: Ronkonkoma, St. James, and Sayville, and their surrounding neighborhoods to test the proposed model. These three stations are treated as candidate mobility hubs. We use the centroids of census tracts within a 5-mile radius of each station as origin nodes, with Manhattan as the destination for a typical weekday morning AM period. Two layers of subnetworks are defined, similar to the toy network. Fig. 3 illustrates the MaaS subnetwork. Each service node is directly connected via MOD service links, with candidate mobility hubs established at each LIRR station. In total, 78 service nodes are included in the network with 113 direct microtransit service links. Additionally, three LIRR links are included to represent line-haul services.
The MOD service considered in this case study is a hypothetical new microtransit operation that operates at unit distance fare rate of \$2\\$2 per mile. For the dummy subnetwork, 78 direct service links parallel to the MOD links are added to represent other access services. Another 78 direct links from service nodes to the Manhattan node are also included in the dummy subnetwork. As a result, the whole network has 244 nodes and 469 links serving the 78 OD pairs. By comparison, the expanded Sioux Falls network in Liu and Chow, (2024) has 30 OD pairs, 82 nodes, and 748 links.
The 78 OD pairs, each originating from a service node to Manhattan, have simulated demand drawn from a normal distribution with a mean of 60 and a standard deviation of 20, representing a typical weekday AM period. The simulated total demand is 4,734, with an average being 60.69 per census tract OD pair. Note that this level of demand falls in the range of observation, as Ronkonkoma Station, the largest one, has 5452 free parking spaces for general passengers, and fills up on a typical weekday. Other input elements are summarized in Table 4. $\bar{d}_{l}$ represents the average link length.
For experimental design, we first use a base case that optimizes link price in the upper level to determine the objective value. The following tests are conducted. (1) To quantify the value of a MH at Ronkonkoma, we then remove the Ronkonkoma station and corresponding MOD links and compare the resulting optimal objective value with the base case. (2) Afterwards, we evaluate a scenario where the microtransit provider increases their price from \$2 to \$3 per mile. (3) The platform considers an alternative objective of maximizing ridership (passenger-miles) under a budget subsidy of \$5000. In this case, we use Eq. (14) to solve problem. (4) Finally, the platform also considers setting the subsidy by MH instead of by link to observe the changes in flow.
All cases are run on a device with an Intel(R) i7-13705 processor. We initiate the proposed algorithm with $\rho_{0}$ equals to 500 and terminate the algorithm either finishing 10 iterations or the optimality gap is below 0.1%. Each iteration has a runtime limit of 10 minutes.
Figure 3: LIRR network illustration: (a) map of LIRR from the MTA with the three stations circled in red; (b) overlay of the census tracts within 5 miles of those stations; and (c) the MaaS subnetwork with links connecting each census tract to nearby stations
Table 4: Subnetwork Mode Parameters
Subnetwork Mode $\mathbf{c^{t}_{l}}$ $\mathbf{c^{o}_{l}}$ $\mathbf{\hat{p}_{l}}$ $\mathbf{\bar{d}_{l}}$
MOD MOD 1.00 0.5 2 5
LIRR 0.80 0 N/A 50
Dummy LIRR dummy 0.90 0 N/A 50
Drive 0.95 0 N/A 60
Park-n-Ride 1.20 0 N/A 5
4.2.1 Base case
The proposed algorithm finds a solution with an optimality gap of less than 0.1% in under 300 seconds. The runtime is significantly shorter than other link-based MaaS design models proposed by previous studies (Liu and Chow,, 2024; Yao and Zhang,, 2024) when solving problems with similar scales. In comparison, the Sioux-Fall network-based case described in Liu and Chow, (2024) requires longer than a 4-hour run time to converge. This illustrates the potential of the proposed algorithm in real-world applications.
The final solution has all MOD service links priced at \$2 per unit length, which is the upper limit of the service price (Remark 3). The total revenue from the MOD service is \$15,959, and 34.19% (1619 passengers) of the total OD demand uses the microtransit service. 752 people use service base at Ronkonkoma station, 583 use service base at St.James station, and 401 use service base at Seyville. This solution suggests a MH platform design with subsidy provision is unnecessary as the pricing offered by the existing microtransit operator is satisfactory to travelers. In other words, there is sufficient incentive for the microtransit operator to enter this last mile market by simply providing service in these areas without further incentives from MaaS fare bundling.
Remark 3.
The model can inform policymakers on the need for a subsidy-enabled MH platform design or to simply implement a traditional transit hub.
4.2.2 Quantifying the value of an MH connecting microtransit to Ronkonkoma station
To quantify the added value of establishing a mobility hub at Ronkonkoma for the microtransit service, we apply the same model to a network that removes the Ronkonkoma station MH (the middle hub in Fig. 3) and its MOD links. The Ronkonkoma station still remains in the dummy network. The modified network has 227 nodes and 418 links. 63 MOD service nodes and 63 MOD links remain in the MaaS platform, with only 15 census tracts losing the MOD service access via Ronkonkoma station. We obtain a solution with an optimality gap less than 0.1% in under 190 seconds. The MOD link prices are still \$2 per mile. Without the Ronkonkoma station serving as a MH, the MaaS platform usage drops from 34.19% down to only 21.4% of total travelers via the remaining two MHs. The total revenue drops to \$10,272. Therefore, establishing a MH at Ronkonkoma station with the microtransit operator is worth \$5,687 per AM period.
Remark 4.
Because the lower-level model is a coalitional choice model, the change in objective value for removing a MH captures the social surplus value of the MH.
4.2.3 Increased microtransit price rate
Because the solution obtained from the base case determines all MH based service links to be priced without subsidy, we consider the microtransit service increasing their price to \$3 per mile. This may reflect higher costs for the microtransit service such that they need to adopt a higher price rate policy.
After 195 seconds, we obtain a solution with the optimality gap less than 0.1%. This time, not all link services are priced at \$3. Instead, 25 links are priced between \$2 to \$3 per unit distance, suggesting that a subsidy is now provided for those links with less than \$3 rates. Fig. 4 illustrates the price link distribution across the network. The farthest nodes from the MHs receive lower prices, i.e. higher subsidies. This is reasonable because the longer the link, the higher the total cost. To attract more customers and generate more revenue, higher subsidies would be the most effective for longer links.
Nevertheless, the high price tolerance of the defined travelers in the area can still be identified. The total revenue is \$20195.8, a 26.5% increase from the base scenario. A total of 30% of OD flow uses the service links, which dropped from 34.19%. The extra revenue of \$4236 from the price change is only feasible with a subsidy-enabled MH design, contrary to the base scenario.
Remark 5.
There exists a price threshold above which a subsidy-enabled MH would outperform a traditional transit hub design.
Figure 4: Changes in link price (left) and flow (right) with increase in price cap from \$2 to \$3.
4.2.4 Flow maximization with link specific subsidy
Alternatively, the LIRR may want to consider an objective of maximizing flow instead of the base scenario. We evaluate this objective with a maximum overall subsidy of \$5000 per AM period. Eq. (14) is used to solve the problem. The algorithm terminates at the first iteration and takes 90 seconds to find a solution with the optimality gap lower that 0.1%. The final penalty term is less than $10^{-6}$, indicating a high accuracy in meeting the lower-level KKT conditions. Fig. 5 illustrates the final link prices and microtransit flows across the network. The total platform reveue is \$17,857.36\\$17,857.36. The revenue minus the \$5000 subsidy leads to a lower net revenue than the base case (by \$3101.64), but ridership increases from 34.19% (1619 passengers) up to 36.70% (1737 passengers), an increase of 118. This makes sense, as the base case shows that subsidy is not needed there, so the trade-off here is simply injecting subsidy to increase ridership at a cost to the platform.
Service areas that are further from the mobility hubs are more likely to receive subsidies, which can be observed from Fig. 5. Longer service link results in higher total service pricing, leading travelers less likely to use MOD services. Therefore, heavy subsidies are focused on further out travelers. For travelers that are closer to the hubs, they are more likely to choose hub based services even they do not receive subsidies.
Figure 5: MOD link price (left) and flow (right)
4.2.5 Maximizing flow with MH-based subsidy
We then enforce a subsidy scheme to allocate the \$5,000 subsidy uniformly for each MH. For all MOD links leading to the same mobility hub, a single subsidy value is applied. The scenario is more constrained, but reflects an alternative subsidy scheme that is dependent on the hub and more practical to control.
The proposed algorithm terminates in 45 seconds when we set the optimality target to be less than 1%, and the penalty term is also less than $10^{-6}$. However, when we set the optimality gap more aggressively to 0.1%, the algorithm fails to reach the goal before 10 iterations. Therefore, it is more computationally demanding to solve the problem using a uniform subsidy scheme. The reason is because of the increased sensitivity towards the subsidy level. When an MH based subsidy value changes, flows on all related MOD links are subject to adjust, which slows down the convergence rate. Therefore, we show the result using the one with an optimality gap of 0.63%.
Table. 5 summarizes the MH based subsidies, which are easier to compare and control than having differing levels of subsidy on every in-platform service link. Clearly the model suggests that trips to Ronkonkoma station should be subsidized 48% more than trips to St. James, likely because of the centrality of Ronkonkoma. The total MOD service link flow is 36.5% (1729 passengers). This is only a 0.5% drop compared with the link-specific subsidy allocation, but the revenue drops by 3% to \$17,288. The loss in revenue and ridership is the trade-off for having a subsidy design that is more convenient to control. Fig. 6 illustrates the differences in price and flow compared to the link-based subsidy result. The spatial distribution of subsidy differences also highlights the imbalance of subsidies with a focus towards more remote areas.
Remark 6.
MH-based subsidies are computationally more expensive than link-based subsidies but provide a subsidy design that is easier to compare and control.
Mobility hub (location in area) Subsidy level (\$/mile)
St. James (top) 0.45
Ronkonkoma (middle) 0.67
Sayville (bottom) 0.61
Table 5: Mobility hub based subsidy
Figure 6: Price difference (left) and flow difference (right) between MH-based and link-based subsidies.
5 Conclusion
In this study, we propose a MH platform design model addressing both the theoretical and computational gaps identified in recent literature for both MaaS and mobility hub based applications. While previous research has largely emphasized the role of MHs as physical transfer points or focused solely on optimal facility location, our approach extends the function of MHs to include their strategic influence on both operator and traveler decision-making, particularly through the explicit modeling of pricing structures and targeted subsidies. By leveraging the PURC framework within a bilevel optimization structure, we capture the complex interplay between public platforms and private mobility operators, while also ensuring computational scalability necessary for real-world, large-scale applications.
The proposed model captures the joint operational planning between travelers and operators needed for effective MaaS platform deployment. The bilevel framework incorporates a revenue-maximizing upper level, where the platform determines optimal pricing or subsidy allocation strategies, and a convex quadratic lower level, where travelers and operators jointly optimize flows and capacities. We transform this bilevel structure into a single-level problem using KKT conditions that guarantees the existence of a global optimum. An augmented penalty-based algorithm is proposed to improve the computational tractability and reduce runtime.
We demonstrate the model application through two sets of numerical examples. The toy example demonstrates the effectiveness of the proposed model and validates the existence of a global optimum. The model can optimally allocate link specific resources from a revenue maximization perspective. We further test the model on a large network based on LIRR stations and their service areas. A total of 244 nodes and 469 links are involved for the base network with 78 OD pairs. Several insights are obtained through the tests:
  • 1.
    The proposed model is much more scalable compared to similar platform design models in past studies. Similar sized problems require hours of runtime ((Liu and Chow,, 2024; Yao and Zhang,, 2024)) while ours only take minutes when setting the right condition.
  • 2.
    The use of the PURC-based lower level problem allows policymakers to quantify the social surplus value of a MH, or the value of enabling subsidy or regulating the price.
  • 3.
    MH-based subsidies are computationally more expensive than link-based subsidies but provide a subsidy design that is easier to compare and control.
The proposed model can be applied to more complex MaaS service designs involving more operators. With more operators, subsidy schemes based on setting a rate per operator can be done. Heterogeneous travel groups can be considered (e.g. splitting each OD pair into portions by income group).
There are multiple gaps remaining. Congestion effects on out-of-platform links can also be added. The link-additive approach could negate non link-additive factors impacting traveler and operator choices. Competition between platforms can be considered by changing the upper level into a generalized Nash equilibrium. Alternative applications of mobility hubs can be studied using this framework: urban air mobility and freight distribution, for example. The model can be further expanded to three-sided markets to deal with electric charging integration.
Acknowledgments
The project is funded by the National Science Foundation (NSF) CMMI-2423908.